The U.S. National Center for Health Statistics reports that 60% of those with artificial hip and knee joints are female; If 30 of these people are selected at random, find the probability that fewer than 15 are female.
Answer #1

We will use Binomial distribution with n=30 and p=0.60

P(Fever than 15 Females) =P(X<15)=P(X<=14)

We will use Excel function ''=BINOMDIST(14,30,0.6,TRUE)'' it will give probability 0.09705

P(P(Fever than 15 Females) =P(X<15)=0.0971 (in Four Decimal)

Hope this will be helpful. Thanks and God Bless You:)